2AlCl3    =>   2Al    +    3Cl2

If 533.1 grams of AlCl3 decomposed at 37.0 °C and 3.02 atm, what would be the volume of chlorine gas?



1
Expert's answer
2021-03-30T07:54:03-0400

Moles of AlCl3 decomposed;

"moles = \dfrac{mass}{MM}=\dfrac{533.1g}{133.34g/mol}=3.998mol"


The balanced equation for the reaction;


2AlCl3 "\to" 2Al + 3Cl2


Moles of Cl2 produced;

Since the mole ratio of AlCl3 to Cl2 is 2:3, then


Moles of Cl2 "=\dfrac{3}{2}xmoles of AlCl3"


"=\dfrac{3}{2}x3.998mol"


"=5.997mol"


Now, at STP, the molar volume of an ideal gas is 22.4L


Assuming ideal situation;


If 1mole of Cl2 will occupy 22.4L

5.997 moles of Cl2 will occupy?


"=\dfrac{5.997molx22.4L}{1mol} = 134.334L"


This is at STP (T1 = 273K, P1 = 1 atm, V1 = 134.334L)


At the new conditions (T2 = 273 + 37 =310K, P2 = 3.02 atm, V2 = ?)


"\dfrac{P1V1}{T1}=\dfrac{P2V2}{T2}"


"P2 = \dfrac{P1V1T2}{T1} = \dfrac{1atm x 134.334Kx310L}{237K}"


= 152.54L